OpenStudy (zenmo):

Evaluate the integral by making the given substitution. (Use C for the constant of integration.)

OpenStudy (zenmo):

\[\int\limits_{}^{}\frac{ \sec^2(\frac{ 1 }{ x^9 }) }{ x^{10} }dx\]\[ u = \frac{ 1 }{ x^9 }\]

jimthompson5910 (jim_thompson5910):

hint: u = 1/(x^9) = x^(-9) du/dx = -9x^(-10) du = -9x^(-10)*dx -du/9 = x^(-10)dx (1/x^10)dx = -du/9

OpenStudy (zenmo):

\[du= 1 / x^9 = \frac{ x^{-10} }{ -10 }\] \[-10du = \frac{ 1 }{ x^{10} }\] \[\int\limits_{}^{}\sec^2(\frac{ 1 }{ x^9 })(\frac{ 1 }{ x^{10} })dx = \int\limits_{}^{}\sec^2u(-10du)\] \[-10\int\limits_{}^{}\sec^2udu = -10tanu +C = -10\tan(1/x^9)+C\] Well, that is my attempt

jimthompson5910 (jim_thompson5910):

let me know if the hint helps or not

jimthompson5910 (jim_thompson5910):

your answer is close but not quite there

OpenStudy (zenmo):

Yea, got it now with that hint, I didn't do the anti-derivative of x^-9 correctly. The answer is: \[-\frac{ 1 }{ 9 }\tan(\frac{ 1 }{ x^9 })+C\]

OpenStudy (zenmo):

Thanks :)

jimthompson5910 (jim_thompson5910):

yep you nailed it